OpenStudy (anonymous):

Ariella had $35 in her wallet. Her friend owed her x dollars and just paid her back. Ariella now has $50 in her wallet. What equation matches this story? A. 15 + x = 50 B. x − 35 = 50 C. 35 + x = 50 D. x − 50 = 35

undeadknight26 (undeadknight26):

This is C because we know how much Arielle has in her wallet but does not know how much her friend owes her.

OpenStudy (anonymous):

c

Parth (parthkohli):

The money she owed would probably be the amount she had in the end MINUS the amount she had initially. Guys, shh! Let them think for themselves for a minute.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Marsha and Nick and 3 of their friends had a yard sale. They made a total of $215.40. They decided to share the total equally. Which number sentence could be used to find s, the amount of money each person received? A. $215.40 ÷ 3 = s B. $215.40 ÷ 5 = s C. $215.40 − 5 = s D. $215.40 × 5 = s

undeadknight26 (undeadknight26):

Which answer do you think it is?

Parth (parthkohli):

Hmm -- what is 215.40 in three equal parts?

undeadknight26 (undeadknight26):

Key words: 3 friends They decided to share the total equally They made a total of $215.40.

OpenStudy (anonymous):

is it a

undeadknight26 (undeadknight26):

Heres a cup of hot cocoa and right you are. XD

OpenStudy (anonymous):

The answer is C